OpenStudy (waheguru):

first the temperature was -16.7 then 14.7 how much did the temperature change by PLZ ANSWER FAST

OpenStudy (anonymous):

Final minus initial.\[14.7-(-16.7)=31.4\]

OpenStudy (waheguru):

CORRECTION FIRST IT WAS -16.1 THEN IT WAS -14.7 SRY

OpenStudy (dumbcow):

i thought i did this already the difference is 1.4 degrees, it went up 1.4

OpenStudy (waheguru):

YEA BUT WHEN WE SUBTRACT I GET -1.4 WHY?

OpenStudy (waheguru):

HOW DO I SHOW MY WORK FOR THIS QUESTION

OpenStudy (waheguru):

DO I ADD?

OpenStudy (anonymous):

Again, final minus initial, it doesn't matter the signs.\[(-14.7)-(-16.1)=2.4\]

OpenStudy (dumbcow):

doesn't matter for difference since you take absolute value however, if you need the change in temp...take 2nd temp and subtract the first -14.7 -(-16.1) = 1.4

OpenStudy (anonymous):

Negative integers can be tricky

OpenStudy (anonymous):

picture the thermometer -16.7 is 16.7 below zero - so to get to zero temp is raised by 16.7. then from o to 14.7 is another rise of 14.7 so total rise of temperature is 14.7 + 16.7 = 31.4

OpenStudy (anonymous):

the subtractuion is 14.7 - ( - 16.7) = 14.7 + 16.7 = 31.4 :- note a '--' = '+'.

OpenStudy (anonymous):

His original question was a typo it was actually negative 14.7 minus negative 16.1